OpenStudy (anonymous):

The length and width of a rectangle are consecutive even integers. The area of the rectangle is 120 square units. What are the length and width of the rectangle?

myininaya (myininaya):

l=n w=n+2 A=lw=n(n+2)=120 n^2+2n-120=0 (n+12)(n-10)=0 n=-12, n=10 lengths are positive so n=10 l=n=10 w=n+2=10+2=12

OpenStudy (anonymous):

Thank you, once again :D !

myininaya (myininaya):

:)

myininaya (myininaya):

if you want you can click good answer the button is right here ^^^^^^ | | | | | | see the arrows above the delete key